OpenStudy (anonymous):

What is the simplified form of 1 over x minus 2 over x squared minus x ? @SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{1}{x} - \frac{2}{x^2}-x}\) like this ?

OpenStudy (anonymous):

Yes but the-x goes like this: 2/x^2 - x

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{1}{x} - \frac{2}{x^2-x}}\) like this you mean ?

OpenStudy (anonymous):

Yes correct

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{1}{x} - \frac{2}{x^2-x}}\) \(\LARGE\color{blue}{ \frac{1}{x} - \frac{2}{x(x-1)}}\) \(\LARGE\color{blue}{ \frac{1\color{darkgoldenrod}{(x-1)}}{x\color{darkgoldenrod}{(x-1)}} - \frac{2}{x(x-1)}}\) \(\LARGE\color{blue}{ \frac{x-1}{x(x-1)} - \frac{2}{x(x-1)}}\) \(\LARGE\color{blue}{ \frac{x-1-2}{x(x-1)} }\) can you finish from here ?

OpenStudy (anonymous):

Thanks @SolomonZelman

OpenStudy (solomonzelman):

Anytime ;)